### v1

# DeepCloth-ROBP&P: Towards a Robust Robot Deployment for Quasi-Static Pick-and-Place Cloth-Shaping Neural Controllers

Abstract
The fidelity gap between simulation-trained vision-based data-driven cloth neural controllers and real-world operation impedes reliable deployment of methods from simulation into physical trials. Real-world grasping errors, such as misgrasping and multilayer grasping, degrade their performance; additionally some fabrics made of synthetic material also tend to stick to the commonly employed Franka Emika Panda’s original gripper. Different approaches adopted various strategies to resolve these problems, further complicating real-world comparison between state-of-the-art methods. We propose DeepCloth-ROBP&P with a simulation-to-reality transfer strategy Towel-Sim2Real and a cloth grasping protocol to consider and mitigate these grasping errors for robustly deploying quasi-static pick-and-place neural controllers in cloth shaping and demonstrate its generalisibility across different deep-learning methods, fabric contexts and robot platforms. Our approach allows us to compare multiple neural controllers in a real environment for the first time, offering valuable insights to the cloth manipulation community.
I Introduction
Cloth shaping algorithms are hard to compare in real-world experiments. There has been impressive progress in recent years on developing vision-based data-driven controllers for cloth flattening and folding, including both reinforcement learning [1, 2, 3, 4, 5] and imitation learning [6, 7]. While such approaches have reported successful transfer of simulation-trained policies to the real-world individually [8], we are aware of no comparative study of different algorithms on a real robot, or analysis of how algorithms generalise across different materials, grippers, or robot platforms. Although many studies have compared these approaches in simulated environments [9, 10, 11], a comprehensive comparison in a real-world setting remains elusive due to the challenges of replicating experiments and accounting for various grasping errors, such as misgrasping, cloth dropping and grasping multiple-layers [1, 7, 10]. In this paper, we analyse the main obstacles to real-world deployment in the cloth-shaping domain and present a first real-world comparison of four recent cloth-shaping algorithms on a variety of flattening and folding tasks, across different platforms and configurations.
One of the primary challenges in deploying cloth manipulation systems in real-world settings can be attributed to the difference between simulated and physical environments, particularly in the behaviour of the fabric itself and the interaction between the end-effector and the fabric. Two main strategies to address these challenges are: (1) enhancing physics engines to more accurately model cloth behaviour to close the reality-to-simulation (Real2Sim) gap which can be considerably difficult to achieve; and (2) developing algorithms that learn directly in real-world, which can be limited to specific use cases and by the scarcity of available demonstrations [4]. We choose to redesign the pipeline for simulation-based algorithms by integrating grasping uncertainty into simulation environments to create more robust policies and enhancing the reliability of policy generation from devising reliable cloth segmentation, motion profile prediction as well as considering the impact of gripper-tip designs on different fabric types.
We present a holistic assessment of four different types of neural controllers’ real-world performance on shaping towels of various sizes, colours, shapes, materials and thickness. We achieve this by following a novel set of strategies for ROBust real-world ROBotic deployment of Quasi-Static Pick-and-Place cloth-manipulation methods. This approach (ROBP&P) is robust, i.e., agnostic, to robot platform, to cloth properties, and to camera-setup – hand-to-eye or eye-in-hand. We make the following contributions:
1. We propose robust strategies covering improved simulation to hardware evaluation as enablers for robust real-world deployment of vision-based cloth manipulation algorithms. The proposed scheme, Towel-Sim2Real, includes (a) a real-to-sim (Real2Sim) benchmark environment incorporating multi-layer grasping and misgrasping uncertainty, (b) an improved flattening oracle policy and (c) a stable cloth segmentation in the real world. We empirically demonstrate the reliability of these schemes on three existing neural controllers allowing for a real-world comparison.
2. We demonstrate the versatility of our Towel-Sim2Real protocol by successfully applying visuomotor diffusion policy [12] to QSP&P towel shaping for the first time, further validating the protocol’s reliability and its applicability to novel neural controllers.
3. We devise a tweezers-extended gripper and a grasping protocol that resolve most of the misgrasping and cloth attachment issues presented in the Franka Emika Panda’s original gripper; this addresses a long-standing issue when it comes to replicating some of the existing work.
4. We conduct extensive experiments in numerous settings to showcase the system compatibility with Franka Emika Panda and Universal Robots UR3e in both ”eye-in-hand” and ”hand-to-eye” configurations under both ROS1 Noetic and ROS2 Humble settings; this demonstrates that our approach works across different robot arms, hardware, and software.
II Related Work
Numerous methods have been applied to single-gripper quasi-static pick-and-place (QSP&P) towel shaping problems but have not been tested in the same robotic environment. We provide such a comparison in this work.
Deep learning controllers
Several influential methods are based on model predictive control (MPC) including VCD [3] and VSF [1]. PlaNet-ClothPick [5], uses a Recurrent State Space Model with variational inference and a domain-specific planning algorithm for fabric flattening. It samples pick locations on the cloth mask using a learned latent dynamic model and optimises policies based on predicted rewards. Among these MPC-based methods, we chose PlaNet-ClothPick as a representative to demonstrate the effectiveness of our deployment approach.
Foldsformer [10], which is a variant of TimeSformer [13], utilises a space-time attention mechanism to capture the instruction feature of folding demonstrations. Another goal-directed method, FabricFlowNet [7], achieves multi-step fabric folding from canonical states using optical flow; however, it does not support folding from crumpled states and its performance was inferior to Foldsformer [10] which we adopt as a SoTA representative for transformer-based controllers to benchmark folding tasks.
JA-TN [11] is a towel-shaping controller that extends the TransporterNet architecture [14] with joint-probability action inference. It uses convolutional neural networks to directly predict pick and place action heatmaps from images. It has demonstrated different types of folding from flattened and crumpled states in simulated environments but has not been extensively evaluated in a real-world experiment. We test JA-TN in real world as a representative of popular TransporterNet-based controllers.
Vision Processing
Visual processing of cloth images often relies on cloth segmentation, with colour thresholding being a common heuristic. VCD [3] focuses on extracting the point cloud of the cloth using segmentation in the target region. Several methods, including FabricFlowNet [7] and Foldsformer [10], utilise cloth-masked depth images as input to their models — to align depth between real and simulated images, they subtract the difference between the average height of the real and the simulated camera. In contrast, VSF [1] employs a domain-randomised [15] strategy centred around the target fabric colour without relying on cloth segmentation. JA-TN [11] also adopts domain randomisation in simulation but allows for a wider range of fabric colours by employing a “segment-anything model” [16] combined with a heuristic mask selection.
Robot Setup
Franka Emika Panda [17] with its original gripper is mainly used to conduct real trials in QSP&P towel shaping [3, 4, 7, 10]. VSF is a notable exception in using a Da Vinci robot with its pincer-like end-effector [1]. This limited diversity in robotic setup represents a significant gap in the field, which our study addresses by demonstrating skill transferability of our robust quasi-static pick-and-place approach (ROBP&P).
The materials used in the literature are often made of cotton, silk, polyester and polyamide materials with side lengths from 7cm to 45cm, typically using a single cloth per experiment. Our research extends this by incorporating multiple fabric types within the same experiment, addressing another crucial gap in existing studies [3, 11].
The community has explored various camera options to capture fabric features, including Zivid One Plus [1], Azure Kinect camera [3], and most commonly, Intel RealSense D435 [7, 4, 10, 11]. Whilst only VCD adopts a side-view camera, other methods employ top-down settings, where the camera is mounted between 0.7 to 0.9 meters above a foamy table surface. Lee et al. [4], FabricFlowNet and Foldsformer adopt ”eye-in-hand” coordination between the end-effectors and the camera; others like VCD and VSF employ ”hand-to-eye” configuration only. Our research bridges this gap by demonstrating the effectiveness of a single method across multiple camera configurations, further enhancing the versatility and applicability of ROBP&P.
Real-world Performance
Due to these differences, to our knowledge, no comprehensive study exists to examine the real-world performance of the towel-shaping neural controllers. Nevetherless, many individual studies reported their own robotic examinations. The transferred algorithm of VCD achieves 94% coverage for cotton square fabric and 95% for silk square fabric [3]. JA-TN reported a 7/10 success rate for a square and 10/10 for a small rectangular synthetic fabric [11] but with human-in-the-loop pick-and-place execution where the policy is generated by the network through real camera input. Lee et al. [4] examined six folding tasks, achieving perfect success rates for four easier folds but struggling with more complex ones. FabricFlowNet and Foldsformer outperform Lee et al. in corresponding foldings [7, 10]. Notably, FabricFlowNet reaches 89% IoU for one-step folding and 60% IoU for multi-step folding, while Foldsformer achieves around 90% IoU for various foldings and outperforms FabricFlowNet in corresponding tasks. Whilst various methods can be compared in the SoftGym simulation benchmark environment, proper real-world comparisons are lacking where typical reported sim2real system failures stem from misgrasping and grasping multiple layers [10, 7, 3, 11]. By allowing various methods to be more easily deployed in a real-world experiment across different platforms, materials, and camera configurations, we present the first real-world comparison of multiple state of the art methods.
Denoising Diffusion Probabilistic Models
DDPMs [18, 19] are a type of generative model that has gained popularity in the image and video generation [20] and various control tasks. The diffusion process involves two key steps: a forward diffusion process and a learned reverse denoising process using a noise predictor. Janner et al. and Huang et al. [21] investigate the use of diffusion models for planning to infer action trajectories executable in specific environments. Wang et al. [22] employ diffusion models for policy representation and regularisation using state-based observations in the realm of reinforcement learning. Chi et al. [12] applied it to behavioural cloning for effective visuomotor control policies; it has been examined successfully on wide range of robot control tasks in Robomimic [23] and Franka Kitchen [24] as well as in bimanual shirt folding. It is trained on 300 demonstrations through a VR setup, achieves a 15/20 success rate with goal-conditioned low-level position control. Our research introduces a novel application of the diffusion method in QSP&P towel-shaping, further demonstrating the reliability of our real-world deployment strategy, TowelSim2Real. In our approach, we leverage oracle policies in simulation environments to collect demonstrations, eliminating the need for costly VR human demonstrations.
III DeepCloth-ROBP&P
We train the domain specific neural controllers such as PlaNet-ClothPick [5], JA-TN [11], as well as a general visuomotor diffusion policy [12] (Section III-C) with our Towel-Sim2Real protocol (Section III-A). The FoldsFormer model was applied without re-training. Additionally, we employ the proposed grasping protocol (Section III-B) for reliable real-world fabric grasping. We collectively refer to this pipeline as DeepCloth-ROBP&P (see Fig. 2).
III-A Towel-Sim2Real
Grasping multiple layers of an article is a common phenomenon in cloth manipulation. It is exacerbated by the fact that most simulation-trained neural controllers do not consider this factor during learning. We propose a new benchmark environment Real2Sim Towels that incorporates real-world behaviours of the cloth-gripper interaction along with a corresponding improved oracle flattening policy to enhance the performance of the transferred deep policies. We enhance the Rainbow Rectangular Fabrics environment [11] by making the fabric share the same colours on both sides, including misgrasping chances and modifying the gripper behaviour to grasp a cluster of particles around the picker instead of precisely picking the closest cloth particle.
We extend the oracle smoothing policy used by the original JA-TN implementation [11] to provide expert trajectories for the trained models. The original policy (Oracle Towel Smoothing, or OTS) leads to errors in real deployment: it typically chooses to unfold a fold on the top of the cloth and reveal a corner if it is hidden underneath the fabric; this policy only works well if a gripper can pick precisely the target particle of the cloth without grasping multiple layers which is challenging in reality. Additionally, we modify the oracle so that the entire fabric remains within the camera’s field of view after flattening, preventing scenarios where portions of the smoothed cloth become obscured. We call the new policy Real2Sim Smoothing (R2S-S).
III-B Cloth Grasping
The original Franka Emika Panda gripper struggles with fabric attachment and precise grasping, and lacks precision for tasks involving edges and corners (see Table 3). We introduce a simple tweezers-extended gripper (Figure 2) integrated into the 3D-printed base, significantly improving control and accuracy. The new gripper may still encounter misgrasping issues due to the precision of action-space conversion. To mitigate this, we implement two key strategies: (1) positioning the pick position 2-pixels inside the cloth-mask border and (2) aligning the picking angle parallel to the cut-line of the pick-point when on a cloth edge. Notably, we refrain from incorporating depth estimation methods, allowing the gripper to descend as deeply as possible.
III-C Visuomotor Diffusion Policy
To demonstrate the robustness and versatility of our method, we apply our protocol to an algorithm never before used in the QSP&P domain: a the diffusion policy. This novel application serves to validate the reliability and adaptability of Towel-Sim2Real across different algorithmic approaches while exploring the potential for expanding the toolset available for QSP&P tasks in fabric manipulation.
Chi et al. (2023) [12] adopt Denoising Diffusion Probabilistic Models (DDPMs) to approximate the conditional policy distribution , where is a sequence of action for prediction horizon from an arbitrary step ; and, is a sequence of history observation with horizon (). Inference denoises the policies for the prediction horizon and only uses -step actions on the environment. A vision encoder produces the vision features by removing the final fully connected layer. The method can also have an internal state vector that concatenates with the to produce conditional global feature .
While sampling, the sampled noisy actions go through iterations of diffusion process, where a diffusion step , last diffused policies and the global condition are fed into noise prediction network to produce prediction noises for step that is used to denoise from :
where and [19]; square cosine scheduler for proposed in iDDPM [25] is used for better performance. Reversely, the DDMP noise scheduler uses the following formula to add noise to the actions
Hence, for denoising step , the ground truth noisy actions are where . Following, the loss function is
where . The method uses Gaussian distribution to normalise the input images and get the statistics from the dataset during the training.
In order to apply the method in our domain, we disabled the down and up sampling of its 1-D convolutions U-Net [26] noise predictor for and as 1, and we keep its enables it for flattening with as 2 and as 4; and we also discard internal state vector for both flattening and folding domains. We train the policy with standard centre-pivot rotation and flipping augmentation along with vision processing strategy proposed by JA-TN [11].
IV Experiments
We conduct shaping experiments on a dark-grey foamy surface with Realsense D435i cameras — UR3e in ”hand-to-eye” configuration (0.74 m above the surface) and Panda in ”eye-in-hand” configuration (0.57 m above the surface on taking images). We use the MoveIt! library [27] to execute QSP&P trajectory, as is common in this domain [7, 10]. During physical trials, humans must reset the initial states for towel-shaping tasks, especially folding, and they also need to manually set the goal states for each trial for setup evaluation [7, 10, 4, 3]. We examine the deployment of flattening controllers on 6 different fabrics varied regarding colour, size, material and shapes (Figure 2). We tested folding performance on Fabrics A and B which are most suitable for folding due to their softness; each fabric is tested 5 times consecutively for each shaping task.
| Method | PlaNet-ClothPick | JA-TN | Diffusion Policy | Human Policy | ||||||||
| Fabric | NC | NI | SR | NC | NI | SR | NC | NI | SR | NC | NI | SR |
| Real2Sim Towels | 91.3 12.1 | 85.3 20.7 | 46.7% | 93.0 13.8 | 87.2 26.3 | 66.7% | 98.2 6.4 | 97.2 10.3 | 90% | - | - | - |
| A | 45.6 14.1 | 12.9 13.5 | 0/5 | 71.9 24.4 | 62.1 32.2 | 2/5 | 68.5 27.0 | 60.3 32.4 | 2/5 | 92.1 13.6 | 89.2 18.7 | 4/5 |
| B | 75.6 13.0 | 57.4 17.2 | 1/5 | 92.5 15.9 | 88.7 24.4 | 4/5 | 85.2 17.0 | 80.0 22.9 | 3/5 | 99.0 1.5 | 98.5 2.1 | 5/5 |
| C | 90.0 18.9 | 84.3 29.1 | 3/5 | 81.7 26.5 | 72.5 40.8 | 3/5 | 96.6 6.2 | 95.2 7.8 | 3/5 | 99.9 0.2 | 99.8 0.4 | 5/5 |
| D | 91.5 15.6 | 87.2 21.7 | 4/5 | 88.0 25.2 | 82.5 36.1 | 3/5 | 96.9 4.0 | 95.7 5.4 | 4/5 | 99.2 0.5 | 98.9 0.7 | 5/5 |
| E | 68.1 25.1 | 54.3 36.1 | 1/5 | 94.1 12.9 | 91.1 18.9 | 4/5 | 87.8 15.2 | 82.6 20.5 | 3/5 | 99.4 1.3 | 98.7 2.47 | 5/5 |
| F | 72.4 12.8 | 58.5 12.6 | 0/5 | 98.29 2.79 | 97.8 3.7 | 4/5 | 81.7 25.0 | 77.2 31.5 | 3/5 | 99.3 1.1 | 98.7 2.5 | 5/5 |
| Real Total | 73.8 22.0 | 59.1 32.7 | 9/30 | 87.8 20.1 | 82.5 28.9 | 20/30 | 86.1 19.1 | 81.8 24.0 | 18/30 | 98.1 5.8 | 97.3 8.0 | 29/30 |
| Folding Tasks | All Corner Inward | Corners Edge Inward | Digonal Cross | Total | ||||||||||
| Fabric | A | B | A | B | A | B | all | |||||||
[…]
| Foldsformer | 83.2 16.6 | 4/5 | 90.2 5.9 | 4/5 | 83.0 5.7 | 0/5 | 77.2 3.3 | 0/5 | 87.1 1.6 | 5/5 | 89.3 1.3 | 5/5 | 85.0 8.4 | 18/30 |
| JA-TN | 83.7 7.8 | 3/5 | 84.1 8.6 | 3/5 | 84.2 7.4 | 3/5 | 79.8 10.5 | 1/5 | 85.6 9.1 | 4/5 | 83.8 10.1 | 4/5 | 83.5 8.4 | 18/30 |
| Diffusion Policy | 91.0 2.1 | 4/5 | 90.0 1.3 | 5/5 | 89.0 4.8 | 3/5 | 84.6 9.8 | 3/5 | 89.1 5.5 | 5/5 | 85.5 3.9 | 5/5 | 88.2 5.4 | 25/30 |
| Human Policy | 93.3 1.9 | 5/5 | 91.4 3.6 | 5/5 | 85.5 11.1 | 4/5 | 90.1 2.9 | 5/5 | 87.0 1.5 | 5/5 | 85.8 3.0 | 4/5 | 88.9 5.6 | 28/30 |
IV-A Generasibility of Towel-Sim2Real
To demonstrate generasability of our deployment protocol Towel-Sim2Real, we conducted experiments with 4 different neural controllers and 6 different fabrics on 4 different shaping tasks. Each physical flattening trial concluded within a maximum of 20 steps normalised coverage exceeding 95% defined as success. We also examine the neural controllers on all-corner inward folding (4 steps), corners-edge inward folding (4 steps) and diagonal-cross foldings (2 steps) with extra 2 steps to allow controllers to recover from mistakes. Intersection-over-Union metric between current and goal masks is used to evaluate step-wise performance, with 85% considered a success. For each shaping task, we also verify success through human judgement. As a control, we also provide a human baselines with a robot executing actions entered via a pick-and-place human interface.
We train PlaNet-ClothPick, JA-TN and diffusion policies in the newly proposed Real2Sim Towels benchmark environment — PlaNet-ClothPick with 20 thousands transitional steps (instead of its original 1 million) using its specially engineered collection policy [5] and the later two with 2000 trajectories created by our updated oracle. Evaluation for flattening controllers in physical trials typically employs step-wise normalised coverage, normalised improvement (requiring real-time cloth segmentation), as well as operational efficiency [1, 3]. Table I shows that all neural policies can transfer successfully to real world scenarios for Fabrics C and D, but all struggle with flattening Fabric A. Although JA-TN performs worse than diffusion policy in simulation, JA-TN does slightly better in real-world in general. Compared to human performance, all methods still needs improvement especially with really soft fabrics like Fabric A.
On the other hand, assessing folding tasks is more complex due to task variety and the self-occlusive nature of goal states, often necessitating human judgement. Intersection-over-Union (IoU) serves as a reasonable proxy metric for step-wise folding performance in real world [7, 10, 4]. The initial state of the fabric are required to be wholly inside the camera view. We only test on the fabrics with low stiffness, such as Fabric A and B, so they will not flip back to their last state.
We retrain JA-TN and diffusion policies with 1000 oracle-based demonstrations [11] in Real2Sim Towels. We also adopt the off-the-shelf Foldsformer model with its original deployment strategy as the baseline for folding comparison. Table II demonstrates that both JA-TN and Diffusion Policy can deploy successfully in reality with our Towel-Sim2Real protocol. Diffusion approaches human performance whilst JA-TN and Foldsformer perform similarly and slightly worse. Despite performing well in all-corner-inward and diagonal-cross foldings, all methods mainly struggle in more complex corners-edge inward folding.
Tables I and II empirically demonstrate that our Towel-Sim2Real training and deployment protocol applies across various neural controller and towel-shaping tasks. It also shows room for improvement on flattening soft fabrics (Fabric A) and performing corners-edge inward folding tasks.
| Fabric | Gripper | SR (%) | Miss (%) | MTL (%) | CA (%) | CD (%) |
| E | Franka | 77.78 | 22.22 | 66.67 | 0.00 | 0.00 |
| Tweezers | 90.00 | 0.00 | 66.67 | 0.00 | 0.00 | |
| C | Franka | 80.00 | 20.00 | 55.56 | 5.56 | 5.56 |
| Tweezers | 93.75 | 0.00 | 61.11 | 0.00 | 0.00 | |
| B | Franka | 83.33 | 16.67 | 69.44 | 2.78 | 5.56 |
| Tweezers | 92.00 | 8.00 | 59.38 | 0.00 | 0.00 | |
| 2-Fold | Franka | 80.56 | 19.44 | 55.56 | 5.56 | 5.56 |
| Tweezers | 91.67 | 8.33 | 61.11 | 0.00 | 0.00 | |
| 3-Fold | Franka | 77.78 | 22.22 | 69.44 | 2.78 | 5.56 |
| Tweezers | 100.00 | 0.00 | 59.38 | 0.00 | 0.00 |
IV-B Robustness of Towel-Sim2Real
Towel-Sim2Real includes three key components: (1) the Real2Sim Towels environment along with its oracle flattening policy R2S-S, (2) a new cloth segmentation heuristics for masking vision inputs and reliable grasping for fabrics, and (3) the image processing for training and deployment proposed in JA-TN [11]. Figure 2 illustrates the reliability of our cloth segmentation method. To assess the significance of the other two components, we compare 4 neural controllers with depth-only input, and we choose Fabric C and D that all policies did well in Table I. We train JA-TN, OTS in Rainbow Rectangular Fabrics and JA-TN, R2S-S in Real2Sim Towels both using JA-TN’s vision processing which also applied to the Diffusion policy. We train another Diffusion, R&S Height with masked-depth input without special processing deployed by adding the difference of simulated and real camera.
Figure 3 shows that the changes in the R2S-S oracle policy represent a substantial leap towards more robust and practical cloth manipulation in real-world robotic applications. It indicates that the diffusion policy trained and deployed with Towel-Sim2Real shows dominant performance on real fabrics, whilst JA-TN, R2S-S and Diffusion, R&S Height only reaches similar performance in simulation. The policies trained with R2S-S demonstration using JA-TN’s vision processing strategy shows better performance that other policies that uses either simple vision processing strategy or trained with OTS policy.
IV-C Grasping with Tweezers-extended Gripper
We evaluate the tweezers-extended gripper against the default Franka gripper aiming to minimise misgrasping, multilayer grasping, cloth dropping and cloth attachment. We encompass a diverse range of cloth conditions including three cloth types (silk, cotton, and polyester), various cloth states (Flat, 2-folds, 3-folds), and multiple grasping scenarios (4 corners, 4 edges, and 2 random cloth points). Table III indicates that the tweezers perform better compared to the Franka gripper across all pick types, particularly excelling in complex graspings on four edges and corners with over 90% success rate. Misgrasp rates for the tweezers remain consistently below 10%, while the Franka Gripper shows higher rates between 16-22%. Both grippers perform similarly for multilayer handling, but the tweezers completely avoid cloth drops and attachment issues, which occur more frequently with the default gripper. Furthermore, the tweezers maintain consistent performance across all cloth types, with particular excellence in handling delicate fabrics like silk.
V Discussion
Many neural controllers in towel shaping only work in specific cloth characteristics regarding particular colour, thickness, and initial cloth positions in camera view [1, 4, 7]; their generalisability to other fabric contexts and robot platform is uncertain. With our protocol DeepCloth-ROBP&P, we deployed four neural policies, including a diffusion policy, in at least two robotic platforms, including Franka Emika Panda with an ”eye-in-hand” setting and UR3e with a ”hand-to-eye” setting.
Limitation
Soft fabrics with high internal friction pose challenges for flattening methods due to the difficulty in separating different parts of the fabric. Conversely, small and stiff fabrics are problematic for folding with a single end-effector as they tend to flip back easily. For all folding controllers, the initial state of the folding tasks should be completely inside the field of vision. Our current robotic system faces several limitations: (1) it struggles to automatically detect success states accurately; (2) it occasionally produces helical and spiral path trajectories; and (3) it still experiences minor misgrasping issues. Fine-tuning camera extrinsic for precise pixel-to-base conversion also remains tedious and time-consuming. Identifying an appropriate workspace for smooth and stable pick-and-place operations is also tricky, primarily due to the robot’s limited reach and the tendency of planned trajectories to produce erratic joint movements when approaching the robot base. Last but not the least, we find that the early checkpoints of the trained networks deploys better for most models, but the surprising diffusion policy does not overfit to the artefact of the simulation under Towel-Sim2Real.
Future
We plan to apply the proposed method to more difficult tasks, such as flattening and folding different garment types. We want to further investigate the capacity of diffusion policy through making folding agnostic to the cloth’s initial state and achieve folding-from-crumpled in the real world. Integrating depth estimation along with a tactile sensor may help to improve the dexterity of the grasping and substantially reduce the effect of grasping multiple layers of the articles, which may be vital for the successful manipulation of complex garment manipulation.
VI Conclusion
We introduce a set of robust robotic strategies, collectively known as DeepCloth-ROBP&P, that enable effective real-world deployment of pick-and-place cloth-shaping neural controllers. We comprehensively evaluate four different neural controllers’ real-world performance on shaping towels with varying size, colour, shape, material and thickness, including a modified visuomotor diffusion policy. Our method also demonstrates the versatility of the proposed protocol across different platforms. Ultimately, this research contributes valuable insights and practical solutions to robotic cloth manipulation, paving the way for more advanced and reliable systems in real-world cloth manipulation.
VII Acknowledgement
We wish to thank Mr Nicol Thomson at the University of St Andrews for setting up the UR3e robot as well as printing the base and integrating the tweezers into the gripper.